OpenStudy (anonymous):

I need help, Im not very good with fractions and I have an assignment with all fractions. 5/2+1/3=1/x

OpenStudy (anonymous):

multiply every term by 6x, no more fractions!

OpenStudy (anonymous):

what do u mean?

OpenStudy (anonymous):

6x*5/2 + 6x*1/3 = 6x*1/x

OpenStudy (anonymous):

5/2 + 1/3 = 1/x (15+2)/6 = 1/x 17/6 = 1/x 17x = 6 x = 6/17

OpenStudy (anonymous):

now what about 4/5z+9/z=1?

OpenStudy (anonymous):

what would I multiply that by?

OpenStudy (anonymous):

This way you can get rid of the fractions quickly instead of working with them Multiply by the least common multiple of the denominator Which in this case is 5z

OpenStudy (anonymous):

z=1/4?

OpenStudy (anonymous):

4 + 45 = 5z

OpenStudy (anonymous):

49/5 = z

OpenStudy (anonymous):

one more, 6/x+1=8/x

OpenStudy (anonymous):

What is the Least Common Multiple of all the denominators?

OpenStudy (anonymous):

x

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so multiply every term by x

OpenStudy (anonymous):

so.....8x=8x...which equals one

OpenStudy (anonymous):

1x

OpenStudy (anonymous):

im lost!!!

OpenStudy (anonymous):

r uthere